OpenStudy (anonymous):

At time t seconds, a particles distance s(t), in micrometers (um), from a point is given by s(t) = e^t - 1. WHat is the average velocity of the particle from t=2 to t=4?

OpenStudy (anonymous):

The average velocity is given by the expression \[\frac{s(4) - s(2)}{4-2}\]

OpenStudy (anonymous):

Evaluate s(4) and s(2) and solve

OpenStudy (anonymous):

Thank you mebs. would the answer be around 23? and do I use um/s?

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

\[\frac{ (e^{4} -1) - (e^{2}-1) }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ e^{4} -1 -e^{2} +1 }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ e^{4} - e^{2} }{ 2 }\]

OpenStudy (anonymous):

= 23.6 um/s

OpenStudy (anonymous):

=24 um/s

OpenStudy (anonymous):

Correct well done

OpenStudy (anonymous):

thank you thats what I got as well. But, why did you do this e4−1−e2+1?

OpenStudy (anonymous):

I simplified it .

OpenStudy (anonymous):

shouldn't it still be e^4-1 - e^2 -1?

OpenStudy (anonymous):

oh nvm i see

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you have to distribute the - sign to al.

OpenStudy (anonymous):

yw =))